# Behavior Modeling Space Reconstruction for E-Commerce Search

### v1

@@ Abstract.
journalyear: 2025conference: Proceedings of the ACM Web Conference 2025; 28 April – 2 May, 2025; Sydney, Australia
@@ 1. Introduction
E-commerce has revolutionized the retail industry by harnessing the capabilities of the World Wide Web to offer consumers round-the-clock access to a global array of products and services (Jain et al., 2021; Wang et al., 2023a, b; Li et al., 2023c; Wang et al., 2022b, 2023c, 2024). In e-commerce applications, the search functionality serves as a critical determinant, matching and ranking relevant items in response to user queries, thereby driving user engagement (e.g., clicks) (Xu et al., 2024; Zhao et al., 2020, 2018b, 2018a). Consequently, a fundamental objective of search model development is to accurately and comprehensively understand user behaviors (i.e., why users interact with certain presented items instead of others) (Lin et al., 2022; Li et al., 2022; Zheng et al., 2022; Li et al., 2023b). However, in practice, available data often only indicates whether a user interacted with an item, without providing explicit insights into the underlying motivations (Liu et al., 2023d, c; Liang et al., 2023; Lin et al., 2023). Thus, effectively modeling user behaviors remains one of the most significant technical challenges in e-commerce search optimization.
@@
Figure 1. Causal graph and Venn diagram.
Table 1. Area indicators for Venn diagram.
Area# 0 1 2 3 4 5
P 0 1 1 1 0 0
R 0 0 0 1 1 1
B 0 0 1 1 1 0
@@ 3.1. Overview
Figure 2 (a) presents the basic structure of the joint modeling frameworks introduced in Section 2.1. The joint model first obtains the query and item representations, $\bm{q}$ and $\bm{r}$, using the text encoder and feature encoder, along with the user representation $\bm{u}$ from the sequence encoder and feature encoder. Subsequently, relevance and preference effects are predicted based on these representations, as described in Equation (1) and Equation (2). These predictions are then merged to generate the behavior prediction, as shown in Equation (3). The proposed DRP operates on the last layer representation of the preference model (within the orange dashed block) and the scoring formula for behavior prediction (within the green dashed block), without constraining the underlying feature encoders or backbone models. Specifically, we design a preference editing framework to eliminate the relevance influence from the preference prediction, as illustrated in Figure 2 (b). We use the last layer representations $(\bm{e}_{p},\bm{e}_{r})$ as neural abstractions for corresponding causal notions $(P,R)$, and identify an orthogonal low-rank matrix $\bm{O}$ that best matches the influence space of $R\to P$. By editing the preference representation in this space and converting the result back to the original space using $\bm{O}^{T}$, we can effectively predict user behavior, capturing both direct relevance and preference effects, as shown by the modified causal graph transitioning from the top left to the top right in the figure. The disentanglement of relevance and preference further allows us to reconstruct a fine-grained modeling space, as depicted in Figure 2 (c). Here, we correctly segment the modeling space into six areas and separately learn their distinct patterns using dual-level adaptive fusion.
@@ 3.4. Inference & Optimization
4. Experiment
4.1. Experiment Setting
4.1.1. Dataset
4.1.2. Backbones & Baselines
4.1.3. Evaluation Metrics
4.1.4. Implementation Details
@@ 4.2. Overall Performance
The existing adaptive fusion method, PRINT, fails to consistently improve joint modeling performance. It divides the relevance effect into three segments: positive, zero, and negative effects, corresponding to Area#4&5, Area#0&3, and Area#1&2, respectively. PRINT struggles to discern the most subtle patterns within combined areas (e.g., Area#1 in Area#1&2), leading to suboptimal outcomes, which are also affected by the unstable Gumbel-Softmax used for area prediction. This calls for a more fine-grained adaptive fusion framework that provides stable predictions across areas.
The proposed DRP outperforms others by incorporating preference editing and adaptive fusion, effectively addressing the issue of entangled effects and reconstructing a fine-grained modeling space. Our work delves into the nuances of the joint modeling problem and avoids the need for human-labeled relevance data.
4.3. Ablation Study
We investigate core components of DRP in this section, focusing on four designed variants to study the effects of orthogonal preference editing and dual-level adaptive scoring. This variant drops the orthogonal constraint on $\bm{O}$. Specifically, it replaces Equation (6) with $\bm{e}_{p_{c}}=\bm{W}_{2}^{T}(\bm{W}_{1}\bm{e}_{p}-\bm{W}_{1}\bm{e}_{r})$, where $\bm{W}_{1}$ and $\bm{W}_{2}$ are learnable matrices with the same dimensions as $\bm{O}$. DRP-2: This variant omits the entire adaptive fusion component to isolate the effectiveness of the remaining component, namely, preference editing. DRP-3: This variant excludes global adaptive fusion to assess its impact. In other words, $\hat{y}_{g}$ in Equation (10) is replaced by $\hat{y}$ from Equation (7). DRP-4: This variant skips Equation (10) to evaluate its efficacy, focusing on the effectiveness of local adaptive scoring.
We evaluate these variants on JDSearch with MLP. The result is listed in Table 4, where AUC and NDCG are presented. From this table, we can get the following conclusions. DRP-1 fails to yield valid results (omitted in the table). Without the orthogonal constraint, the editing space learned by $\bm{W}_{1}$ may not align with the causal intervention space of relevance on preference, potentially subtracting predictive information about user behaviors. Moreover, the learned reverse transformation $\bm{W}_{2}$ is not necessarily the strict inverse of $\bm{W}_{1}$, which could result in missing the original prediction space captured by $\bm{W}_{p}$ in Equation (7). DRP-2 still outperforms the ‘Base’, suggesting that even with just preference editing, joint modeling can benefit from disentangled relevance and preference effects. DRP-3 and DRP-4 perform better than DRP-2, underscoring the effectiveness of global and local adaptive fusion, respectively.
@@ 4.5. Model Visualization
This section visualizes the effects of dual-level adaptive fusion described in Section 3.3. We load the checkpoint of DRP with MLP and HEM on KuaiSAR, subsequently calculating the average behavior predictions for each designated area (Area#0-5), illustrated via a heatmap in Figure 4. We classify samples with the top 20% relevance scores as relevant and the same for preference to construct areas. For illustrative clarity, we incorporate an additional Area#0 following Area#6. The modeled areas depicted in Figure 4 (I-III) correspond directly to the configuration shown in Figure 2(c). From this figure, several observations emerge: (i) The fixed fusion in Equation 7 fails to differentiate Area#2&4 and Area#1&5, collapsed as Area#3. (ii) Through global adaptive fusion, DRP successfully discriminates among four distinct areas: Area#0, Area#1&2, Area#3, and Area#4&5 (noted by similar color in the figure). (iii) The implementation of local adaptive fusion ultimately enables discerning fine-grained patterns across various areas. These findings are in accord with our theoretical analysis in Section 3.3, underscoring the rationale behind our design.
@@ 5. Related Works
E-commerce search models, also known as product search models, can be divided into three categories: ad-hoc search models, personalized product search, and the development of relevance models. Ad-hoc models (Robertson et al., 2009; Huang et al., 2013; Ai et al., 2017) conceptualize e-commerce search as a user-independent relevance modeling problem. While personalized search endeavors to synthesize user behavior sequences to generate personalized results, which is enhanced through the utilization of knowledge graphs (Ai and Narayanan. R, 2021) and transformers (Ai et al., 2019a; Bi et al., 2020). Xiao et al. (2019) was the first to articulate the notion of relevance modeling, which has inspired numerous subsequent efforts aimed at designing pre-trained relevance models to be integrated within a preference modeling framework, operating as a joint behavior modeling approach. Techniques such as knowledge distillation (Jiang et al., 2020), self-supervised learning (Chen et al., 2023), and pretraining pipelines (Zan et al., 2023) are employed in this context. Nonetheless, these frameworks currently exhibit a heavy reliance on laboriously collected relevance data; although some enhancement methods based on click data have been proposed (Yao et al., 2021), there remains unexplored for the joint modeling framework to address latent issues or augment its applicability. A relevant contribution might be PRINT (Hong et al., 2024), which personally merges the relevance effect during the fusion phase in a vague manner. Our study examines the challenges posed by the collapsed modeling space problem within the joint framework, and we introduce DRP, which employs fine-grained, explicit adaptive fusion for easy deployment across diverse systems without relevance data. Another relevant research area is the solution for the collapsed modeling space. Despite substantial attention from CVR prediction studies (Zhu et al., 2023; Huang et al., 2024; Wang et al., 2022a) and approaches addressing impression bias (Lin et al., 2024; Lee et al., 2022), joint modeling strategies remain limited explorations. We are pioneering a thorough analysis of the collapsed modeling space issue within the context of relevance-preference joint modeling.

### v2

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3696410.3714949isbn: 979-8-4007-1274-6/25/04ccs: Information systems Information retrieval
@@ 1. Introduction
E-commerce has revolutionized the retail industry by harnessing the capabilities of the World Wide Web to offer consumers round-the-clock access to a global array of products and services (Jain et al., 2021; Wang et al., 2023a, b; Li et al., 2023c; Wang et al., 2022c, 2023c, 2024). In e-commerce applications, the search functionality serves as a critical determinant, matching and ranking relevant items in response to user queries, thereby driving user engagement (e.g., clicks) (Xu et al., 2024; Zhao et al., 2020, 2018b, 2018a). Consequently, a fundamental objective of search model development is to accurately and comprehensively understand user behaviors (i.e., why users interact with certain presented items instead of others) (Lin et al., 2022; Li et al., 2022; Zheng et al., 2022; Li et al., 2023b). However, in practice, available data often only indicates whether a user interacted with an item, without providing explicit insights into the underlying motivations (Liu et al., 2023d, c; Liang et al., 2023; Lin et al., 2023). Thus, effectively modeling user behaviors remains one of the most significant technical challenges in e-commerce search optimization.
@@ 2.1. Joint Modeling Framework
Figure 1. Causal graph and Venn diagram.
Table 1. Area indicators for Venn diagram.
Area# 0 1 2 3 4 5
P 0 1 1 1 0 0
R 0 0 0 1 1 1
B 0 0 1 1 1 0
@@ 3.1. Overview
Figure 2 (a) presents the basic structure of the joint modeling frameworks introduced in Section 2.1. The joint model first obtains the query and item representations, $\bm{q}$ and $\bm{r}$, using the text encoder and feature encoder, along with the user representation $\bm{u}$ from the sequence encoder and feature encoder. Subsequently, relevance and preference effects are predicted based on these representations, as described in Equation (1) and Equation (2). These predictions are then merged to generate the behavior prediction, as shown in Equation (3). The proposed DRP operates on the last layer representation of the preference model (within the orange dashed block) and the scoring formula for behavior prediction (within the green dashed block), without constraining the underlying feature encoders or backbone models. Specifically, we design a preference editing framework to eliminate the relevance influence from the preference prediction, as illustrated in Figure 2 (b). We use the last layer representations $(\bm{e}_{p},\bm{e}_{r})$ as neural abstractions for corresponding causal notions $(P,R)$, and identify an orthogonal low-rank matrix $\bm{O}$ that best matches the influence space of $R\to P$. By editing the preference representation in this space and converting the result back to the original space using $\bm{O}^{T}$, we can effectively predict user behavior, capturing both direct relevance and preference effects, as shown by the modified causal graph transitioning from the top left to the top right in the figure. The disentanglement of relevance and preference further allows us to reconstruct a fine-grained modeling space, as depicted in Figure 2 (c).
@@ 3.4. Inference & Optimization
4. Experiment
We present experimental settings and extensive empirical results in this section. Additional experiments are provided in Appendix A.
4.1. Experiment Setting
4.1.1. Dataset
4.1.2. Backbones & Baselines
4.1.3. Evaluation Metrics
4.1.4. Implementation Details
@@ 4.2. Overall Performance
The existing adaptive fusion method, PRINT, fails to consistently improve joint modeling performance. It divides the relevance effect into three segments: positive, zero, and negative effects, corresponding to Area#4&5, Area#0&3, and Area#1&2, respectively. Thus, it struggles to discern subtle patterns within combined areas (e.g., Area#1 in Area#1&2), leading to suboptimal outcomes. This calls for a more fine-grained adaptive fusion framework that provides stable predictions across areas.
The proposed DRP outperforms others by incorporating preference editing and adaptive fusion, effectively addressing the issue of entangled effects and reconstructing the correct modeling space. Our work delves into the nuances of the joint modeling problem and avoids the need for human-labeled relevance data.
@@
DRP-5 0.6813 0.2045 0.6876 0.2047 0.6851 0.2060
4.3. Ablation Study
We investigate core components of DRP in this section, focusing on four designed variants to study the effects of orthogonal preference editing and dual-level adaptive scoring.
  • DRP-1 drops the orthogonal constraint on $\bm{O}$. Specifically, it replaces Equation (6) with $\bm{e}_{p_{c}}=\bm{W}_{2}^{T}(\bm{W}_{1}\bm{e}_{p}-\bm{W}_{1}\bm{e}_{r})$, where $\bm{W}_{1}$ and $\bm{W}_{2}$ are learnable matrices with the same dimensions as $\bm{O}$.
  • DRP-2 omits the adaptive fusion component to isolate the effectiveness of the remaining component, i.e., preference editing.
  • DRP-3 excludes global adaptive fusion to assess its impact, i.e., $\hat{y}_{g}$ in Equation (10) is replaced by $\hat{y}$ from Equation (7).
  • DRP-4 skips Equation (10) to evaluate its efficacy, focusing on the effectiveness of local adaptive scoring.
  • DRP-5 omits the preference editing module and directly constructs the modeling space using the impure preference effect.
We evaluate these variants on JDSearch with MLP. The result is listed in Table 4, where AUC and NDCG are presented. From this table, we can find the original design achieves the best results, outperforming all variants. This suggests that the combination of adaptive fusion with disentangled effects greatly enhances effectiveness, highlighting the success of the reconstructed modeling space by DRP.
@@ 4.5. Model Visualization
This section visualizes the effects of dual-level adaptive fusion described in Section 3.3. We load the checkpoint of DRP with MLP and HEM on KuaiSAR, subsequently calculating the average behavior predictions for each designated area (Area#0-5), illustrated via a heatmap in Figure 4. We classify samples with the top 20% relevance scores as relevant and the same for preference to construct areas. For illustrative clarity, we incorporate an additional Area#0 following Area#6. The modeled areas depicted in Figure 4 (I-III) correspond directly to the configuration shown in Figure 2(c). From this figure, several observations emerge: (i) The fixed fusion in Equation 7 fails to differentiate Area#2&4 and Area#1&5, collapsed as Area#3. (ii) Through global adaptive fusion, DRP successfully discriminates among four distinct areas: Area#0, Area#1&2, Area#3, and Area#4&5 (noted by similar color in the figure). (iii) The implementation of local adaptive fusion ultimately enables discerning fine-grained patterns across various areas. These findings are in accord with our theoretical analysis in Section 3.3.
@@ 5. Related Works
E-commerce search models, also known as product search models, can be divided into three categories: ad-hoc search models, personalized product search, and the development of relevance models. Ad-hoc models (Robertson et al., 2009; Huang et al., 2013; Ai et al., 2017) conceptualize e-commerce search as a user-independent relevance modeling problem. While personalized search endeavors to synthesize user behavior sequences to generate personalized results, which is enhanced through the utilization of knowledge graphs (Ai and Narayanan. R, 2021) and transformers (Ai et al., 2019a; Bi et al., 2020). Xiao et al. (2019) was the first to articulate the notion of relevance modeling, which has inspired numerous subsequent efforts aimed at designing pre-trained relevance models to be integrated within a preference modeling framework, operating as a joint behavior modeling approach. Techniques such as knowledge distillation (Jiang et al., 2020), self-supervised learning (Chen et al., 2023), and pretraining pipelines (Zan et al., 2023) are employed in this context. Nonetheless, these frameworks currently exhibit a heavy reliance on laboriously collected relevance data; although some enhancement methods based on click data have been proposed (Yao et al., 2021), there remains unexplored for the joint modeling framework to address latent issues or augment its applicability. A relevant contribution might be PRINT (Hong et al., 2024), which personally merges the relevance effect during the fusion phase in a vague manner. Our study examines the challenges posed by the collapsed modeling space problem within the joint framework, and we introduce DRP, which employs fine-grained, explicit adaptive fusion for easy deployment across diverse systems without relevance data. Another relevant research area is the solution for the collapsed modeling space. Despite substantial attention from CVR prediction (Zhu et al., 2023; Huang et al., 2024; Wang et al., 2022a) and impression bias (Lin et al., 2024; Lee et al., 2022), joint modeling strategies remain limited explorations. We are pioneering a thorough analysis of the collapsed modeling space within the context of relevance-preference joint modeling.
@@ Acknowledgements.
Appendix A Additional Results
A.1. Comparison on DeepFM & DESTINE
Table 5. Additional overall comparison on Private.
PM RM JM AUC Logloss NDCG HR
DeepFM DSSM Base 0.6619 0.05606 0.1581 0.2820
DRP 0.6631 0.05578 0.1623 0.2867
QEM Base 0.6658 0.05575 0.1635 0.2885
DRP 0.6707 0.05558 0.1658 0.2919
HEM Base 0.6661 0.05574 0.1633 0.2878
DRP 0.6697 0.05562 0.1650 0.2901
DESTINE DSSM Base 0.6709 0.05565 0.1660 0.2921
DRP 0.6715 0.05568 0.1671 0.2936
QEM Base 0.6697 0.05563 0.1657 0.2910
DRP 0.6709 0.05554 0.1668 0.2924
HEM Base 0.6706 0.05564 0.1662 0.2919
DRP 0.6711 0.05557 0.1676 0.2938
We present performance of DRP with preference model DeepFM and DESTINE on Private in Table 5. Our findings indicate that DRP outperforms ‘Base’ when using these PM backbones.
A.2. Comparison with Disentanglement Method
Table 6. Additional disentanglement baseline on KuaiSAR.
PM RM JM AUC Logloss NDCG HR
MLP DSSM InvPref 0.6211 0.3554 0.4114 0.7117
DRP 0.6229 0.3549 0.4168 0.7145
QEM InvPref 0.6122 0.3621 0.4101 0.7148
DRP 0.6227 0.3551 0.4143 0.7132
HEM InvPref 0.6218 0.3552 0.4110 0.7121
DRP 0.6230 0.3551 0.4138 0.7122
DCN DSSM InvPref 0.6206 0.3556 0.4131 0.7112
DRP 0.6200 0.3557 0.4145 0.7133
QEM InvPref 0.6200 0.3555 0.4125 0.7134
DRP 0.6243 0.3547 0.4163 0.7142
HEM InvPref 0.6223 0.3550 0.4136 0.7134
DRP 0.6236 0.3545 0.4152 0.7144
We compare the proposed DRP with the latest disentanglement method InvPref (Wang et al., 2022b) on KuaiSAR. The results is presented in Table 6. DRP outperforms InvPref in most scenarios. However, InvPref demonstrates stronger compatibility when compared to other baselines in Table 3.